# Quantum algorithms for viscosity solutions to nonlinear Hamilton-Jacobi equations based on an entropy penalisation method

### v1

@@ Abstract
We present a framework for efficient extraction of the viscosity solutions of nonlinear Hamilton–Jacobi equations with convex Hamiltonians. These viscosity solutions play a central role in areas such as front propagation, mean-field games, optimal control, machine learning, and a direct application to the forced Burgers’ equation. Our method is based on an entropy penalisation method proposed by Gomes and Valdinoci [1], which generalises the Cole–Hopf transform from quadratic to general convex Hamiltonians, allowing a reformulation of viscous Hamilton-Jacobi dynamics by a discrete-time linear dynamics which approximates a linear heat-like parabolic equation, and can also extend to continuous-time dynamics. This makes the method suitable for quantum simulation. The validity of these results hold for arbitrary nonlinearity that correspond to convex Hamiltonians, and for arbitrarily long times, thus obviating a chief obstacle in most quantum algorithms for nonlinear partial differential equations. We provide quantum algorithms – both analog and digital – for extracting pointwise values, gradients, minima, and function evaluations at the minimiser of the viscosity solution, without requiring nonlinear updates or full state reconstruction.
@@ I Introduction
The simulation of partial differential equations (PDEs) on quantum computers has emerged as a promising direction in computational mathematics and quantum information science. Most existing quantum algorithms for PDEs, however, are restricted to linear systems. In contrast, nonlinear PDEs of Hamilton-Jacobi type, which appear ubiquitously in mechanics, control theory, and statistical physics, remain a formidable challenge for quantum computation. The goal of this paper is to develop efficient quantum algorithms for viscosity solutions to the Hamilton-Jacobi equations.
@@
So far, there have been two broad classes of quantum algorithms for nonlinear PDEs. The first one uses analytic transformations to make the equations linear, which we call linear representation [6]. The Level set [3] and von-Neumann Koopman approach [7] belong to this category. The other class of quantum algorithms for nonlinear ODEs and PDEs, for example the Carleman truncation method [8, 9], approximates the original problems by linear ones, which we classify as linear approximations in [3]. For nonlinear PDEs like Hamilton-Jacobi or fluid dynamic equations, such linear approximations fail to capture essential nonlinear physical phenomena such as caustics, shocks and turbulence. Linear representations – if they exist and one can design them – can capture these nonlinear features. Another restriction of the linear approximation methods is that they are usually valid for weak nonlinearity and with strong dissipations. Linear representation methods are expected to be valid for general nonlinearity, and globally in time. However, finding a linear representation is not easy.
@@
In this paper, we use what we call a linear formulation, where we can still approximate the results using a linear representation where the results still remain valid globally in time and truly nonlinear features can be captured. One challenge comes from the numerical viscosity always needed for computing discontinuous solutions in order to eliminate numerical oscillations and guarantee convergence to the physically relevant viscosity solutions. Even for a linear $\mathcal{H}$, the numerical viscosity in modern shock or caustic capturing schemes is nonlinear [10]. To avoid this difficulty, we use a linear, artificial viscosity approximation, which perturbs the Hamilton-Jacobi equation into a viscous Hamilton-Jacobi equation with a linear viscosity term. The concept of artificial viscosity was first introduced by von Neumann and Ritchmyer [11] in the 1950’s for the computation of shock waves in compressible gas dynamics, and has become a standard numerical tool to handle singular solutions. To overcome the problem of nonlinear $\mathcal{H}$, we first observe that for quadratic Hamiltonians, the classical Cole–Hopf transformation converts the viscous Hamilton-Jacobi equation into a linear heat equation, providing an exact mapping – thus a true linear representation – between nonlinear and linear dynamics. For more general convex Hamiltonians, we adopt and generalise the entropy penalisation method of Gomes and Valdinoci [1], which constructs an entropy-regularised approximation that correspond to the viscous Hamilton-Jacobi equations. Furthermore, via the Cole-Hopf transform, it gives a linear discrete-time model that corresponds to a linear parabolic PDE, which is convenient for quantum simulation. This transformation enables a linear formulation for the nonlinear viscous Hamilton-Jacobi equation without loss of essential nonlinear behavior. A related method was introduced in [12] to compute the minimal Mather measure.
@@ II.1 Background on viscosity solutions of Hamilton-Jacobi equations
Instead, we use a linear, artificial viscosity approximation. The simplest such approximation is the following regularised problem, also called the viscous Hamilton-Jacobi equation
@@
Then by combing Theorem 3.1 and Corollary 3.4 from [23] gives:
@@
$\sup_{0\leq t\leq T}\|\nabla S(t,x)-\nabla S_{\nu}(t,x)\|_{l_{2}(\mathbb{T}^{d})}\leq C{\nu}^{\beta},$ (4)
where $\beta\in(1/2,1)$, $C$ is linear in $d$, scales with $T^{3/2}$, and also depends on $S_{0}$ and $H$.
@@ II.2.1 Quadratic Hamiltonian
$S_{\nu}(t,x)=-2\nu\log(\mathcal{N}_{0}u(t,x)),\qquad\mathcal{N}_{0}=\int e^{-S_{\nu}(0,x)/\nu}dx.$ (6)
@@ II.2.2 More general convex Hamiltonians
In the previous section, the Hamiltonian is quadratic in the canonical momentum, which is represented by the gradient of $S_{\nu}(t,x)$, and intuitively this term acts like a kinetic energy term. For Hamiltonians with terms that are more general in canonical momentum, however, it is not possible to transform the nonlinear PDE into a linear one. So we proceed using an alternative method introduced in [1], which still makes use of the Cole-Hopf transformation, though not directly.
@@
The basic idea proposed in [1] is that, instead of working directly with the PDEs in continuous time and then performing the Cole-Hopf transformation, one first chooses a small time step $h>0$ and begins with discrete-time schemes. Here we define at time-step $n$ the solution $S_{D}^{n}(x)$ via a time-marching process:
@@
where $G$ is a nonlinear operator that approximately corresponds to the viscous Hamilton-Jacobi equation Eq. (2). The solution of the discrete-time scheme $S_{D}^{n}$ can be related to the solution of another discrete-time scheme $u^{n}$, defined by
$u^{n+1}(x)=L[u^{n}](x)=\int e^{-h\mathcal{L}(x,v)/(2\nu)}u^{n}(x+hv)dv,$ (13)
Here $L$ is a linear operator that satisfies the property $\|L[u]\|\leq\|u\|$, hence the dynamics (13) is contractive, and can be shown to be approximation of some linear parabolic heat-like operator when the discrete time step is small. With the above defined $G$ and $L$, the solutions $S_{D}^{n}(x)$ and $u^{n}(x)$ are then related through the Cole-Hopf transformation.
@@
    Solve for $u^{n}$, for $n=0,\cdots,N$, by the linear scheme (13)
@@
    Define $\tilde{u}^{N}(x)=Z_{L}(x)^{N}u^{N}(x)$, where $Z_{L}(x)=\int_{\mathbb{R}^{d}}e^{-h\mathcal{L}(x,v)/(2\nu)}dv$ is the normalization constant;
  4. (4)
    The solution to $S_{D}^{N}(x)$ can be recovered by applying the Cole-Hopf transformation on $\tilde{u}^{N}(x)$
    $S_{D}^{N}(x)=-2\nu\log\tilde{u}^{N}(x),$ (14)
    where $S_{D}^{n}(x)$ is the solution to the discrete scheme (12);
  5. (5)
    The solution of the viscous Hamilton-Jacobi equation (1) at time-step NN is recovered by
    $S^{N}_{\nu}=S_{D}^{N}(x)+2\nu\,\ln Z_{L}(x);$ (15)
    where $Z_{L}(x)=\int_{\mathbb{R}^{d}}e^{-hL(x,v)/(2\nu)}dv$;
@@
    $\nabla S_{\nu}^{N}(x)=-2\nu\nabla\tilde{u}^{N}(x)/\tilde{u}^{N}(x).$ (16)
@@
where an entropy regulariser $\gamma(v)\ln\gamma(v)$ is added for $\nu\neq 0$ and $\gamma(v)$ is a probability density over $\mathbb{R}^{d}$ so $\int_{\mathbb{R}^{d}}\gamma(v)dv=1$. The inclusion of an entropy regularisation [25] introduces uncertainty in $v$ and the optimal probability distribution $\gamma(v)$ is expected to take the form of a Gibbs state with temperature $\nu$, from well-known arguments related to free energy minimisation. Eq. (17) is also a Hopf-Lax-like representation, one that uses an entropy regulariser. Using this definition, one can suggestively see how the Hamilton-Jacobi equation emerges. For small $\nu$ and $h$, using Eq. (17), the time derivative of $S_{D}$ behaves roughly like $\partial S_{D}^{n}/\partial t\sim(S_{D}^{n+1}-S_{D}^{n})/h=(G[S_{D}^{n}]-S_{D}^{n})/h\sim\inf(\mathcal{L}-p\cdot v)=-\mathcal{H}(x,p)$, where the $p\cdot v$ term comes from the Taylor expansion $S_{D}^{n}(x+hv)\sim S_{D}^{n}(x)+h\nabla S_{D}^{n}\cdot v$ and $p=-\nabla S_{D}$. This then gives the Hamilton-Jacobi equation $\partial S_{D}/\partial t+\mathcal{H}(x,p)=0$. This is to provide intuition only, and for full details see [1].
@@
The steps in the algorithm above is in discrete-time. However, a continuous-time limit (setting $h\rightarrow 0$), which we see later is more suitable for the quantum simulation of PDEs. To derive the continuous-time limit, it is first necessary to ensure that the following linear iterative process is stable
@@
$\frac{\partial u(t,x)}{\partial t}=\sum_{i=1}^{d}\mu_{i}\frac{\partial u(t,x)}{\partial x_{i}}+\sum_{j,k=1}^{d}\nu_{jk}\frac{\partial^{2}u(t,x)}{\partial x_{j}\partial x_{k}},\qquad u(0,x)=\frac{e^{-S(0,x)/(2\nu)}}{\mathcal{N}_{0}},\qquad\mathcal{N}_{0}=\int e^{-S(0,x)/\nu}dx,$
$\mu_{i}=v_{i}^{*},\qquad\nu_{jk}=v^{*}_{j}v^{*}_{k}+2\nu D^{-1}_{jk},\qquad v^{*}=\text{argmin}_{v}K(v),\qquad D_{jk}=\partial^{2}K(v^{*})/\partial v_{j}\partial v_{k},$ (25)
where the drift coefficient $\mu_{i}$ and the diffusion coeffients $\nu_{ij}$ are all independent of $x$. Let $S_{\nu}(t,x)$ be the solution of the viscous Hamilton-Jacobi equation in Eq. (2) with
@@
    Solve for $u(t,x)$ using the linear parabolic PDE in Eq. (5) in Lemma 5
@@
    Apply the Cole-Hopf transformation to $u(t,x)$ to obtain $S_{\nu}(t,x)=-2\nu\log u(t,x)$.
@@ III.2 Digital quantum simulation
If we instead embed the solution of our problem into a discrete-variable quantum state $|u(t)\rangle_{DV}$, then we need to first decompose $x\in[-1/2,1/2]^{d}$ into a grid of size $N_{x}$ in each dimension, and consider the cost of simulating the state of $d\log N_{x}$ qubits (or a $N_{x}^{d}$-dimensional qudit), that embeds the solution of Eq. (27)
@@
Each $j_{k}$, $k=1,\cdots d$ can be decomposed into its binary decimal representation $j_{k}=j_{k}^{(1)}\cdots j_{k}^{(\log N_{x})}$, where each $j_{k}^{(l)}\in\{0,1\}$.
@@
which is a matrix of size $N_{\eta}N_{x}^{d}\times N_{\eta}N_{x}^{d}$ that acts on a system of $\log N_{\eta}+d\log N_{x}$ qubits. One can obtain $\mathbf{A}_{1,DV}$ and $\mathbf{A}_{2,DV}$ by discretising the spatial components of $\mathbf{A}_{1},\mathbf{A}_{2}$, for example. In this case replace $\hat{p}\rightarrow\hat{P}$, $\hat{x}\rightarrow{X}$, where $\hat{P}_{j}$ is a sparse shift matrix and $\hat{X}=\text{diag}(-1/2N_{x},\cdots,1/2N_{x})$ for the one-dimensional case, and $\hat{D}=\text{diag}(-1/2N_{\eta},\cdots,1/2N_{\eta})$. We summarize the algorithm below.
@@
    Apply projective operator $\hat{\Pi}_{\xi>0}$ on $\log N_{\xi}$ ancilla qubits (or $N_{\xi}$-dimensional ancilla qudit). Flag success – output $|u(t)\rangle$ – when $\xi>0$ and success probability $p_{succ}=\|\mathbf{u}(t)\|/\|\mathbf{u}(0)\|$ (amplitude amplification used for quadratic boosting).
@@
Let $S_{\nu,\Delta}=-2\nu\log|u\rangle$ where $|u\rangle$ is the discrete solution obtained by Algorithm 4. Then,
@@
$\tilde{O}\left(\frac{1}{\|\mathbf{u}(t)\|}\left(\alpha_{A}t\mu+\log\left(\frac{\mu}{\epsilon\|\mathbf{u}(t)\|}\right)\right)\right),\qquad\|\mathbf{u}(0)\|=1,$ (39)
where $\alpha_{A}\geq\|\mathbf{A}_{1}\|,\|\mathbf{A}_{2}\|$ (spectral norm) and using $O(\|\mathbf{u}(0)\|/\|\mathbf{u}(t)\|)$ queries to the initial state preparation protocol for $|v(0)\rangle$. If one discretises the initial ancilla state $|\Xi\rangle_{DV}$ by discretising $\xi$ with the total number of grids $N_{\xi}\sim 1/\epsilon$, then $\mu=O(1/\epsilon)$. There also exists optimal choices of the ancilla initial state that can give $\mu=O(\log(1/\epsilon))$. For time-dependent Hamiltonians, the optimal query complexity only incurs an extra $\log(1/\epsilon)$ factor.
@@
Ignoring the terms with logarithmic scaling, one can simulate $|u(t)\rangle_{DV}$ to precision $\epsilon$ with cost $\tilde{O}(\alpha_{A}t\log(1/\epsilon)/\|\mathbf{u}(t)\|)$, where it is sufficient to choose $\alpha_{A}\geq\max\{\|\mathbf{A}_{1}\|,\|\mathbf{A}_{2}\|\}=O(d^{4}/\epsilon^{2})$.

### v2

@@ Abstract
We present a framework for efficient extraction of the viscosity solutions of nonlinear Hamilton–Jacobi equations with convex Hamiltonians. These viscosity solutions play a central role in areas such as front propagation, mean-field games, optimal control, machine learning, and a direct application to the forced Burgers’ equation. Our method is based on an entropy penalisation method proposed by Gomes and Valdinoci [1], which generalises the Cole–Hopf transform from quadratic to general convex Hamiltonians, allowing an approximation of viscous Hamilton-Jacobi dynamics by a discrete-time linear dynamics which approximates a linear heat-like parabolic equation, and can also extend to continuous-time dynamics. This makes the method suitable for quantum simulation. The validity of these results hold for arbitrary nonlinearity that correspond to convex Hamiltonians, and for arbitrarily long times, thus obviating a chief obstacle in most quantum algorithms for nonlinear partial differential equations. We provide quantum algorithms – both analog and digital – for extracting pointwise values, gradients, minima, and function evaluations at the minimiser of the viscosity solution, without requiring nonlinear updates or full state reconstruction.
@@ I Introduction
The simulation of partial differential equations (PDEs) on quantum computers has emerged as a promising direction in computational mathematics and quantum information science. Most existing quantum algorithms for PDEs, however, are restricted to linear systems. Quantum simulation of nonlinear PDEs remains a formidable challenge. In this paper we develop efficient quantum algorithms for viscosity solutions to nonlinear PDEs of Hamilton-Jacobi type, which appear ubiquitously in mechanics, control theory, and statistical physics.
@@
So far, there have been two broad classes of quantum algorithms for nonlinear PDEs. The first one uses analytic transformations to make the equations linear, which we call linear representation [6]. The level set [3] and von-Neumann Koopman approach [7] belong to this category. Such methods have no restrictions for the strength of nonlinearity but are limited in its applicability to just few nonlinear PDEs. The other class of quantum algorithms for nonlinear ODEs and PDEs, for example the Carleman truncation method [8, 9], approximates the original problems by linear ones, which we classify as linear approximations in [3]. For nonlinear PDEs like Hamilton-Jacobi or fluid dynamic equations, such linear approximations are unable to capture essential nonlinear physical phenomena such as caustics, shocks and turbulence, since they appear in physical systems with strong nonlinearity and weak dissipation. In fact a main restriction of the linear approximation methods is that they are usually valid for weak nonlinearity and with strong dissipations. The nonlinear Hamilton-Jacobi equation, in the contrary, has strong nonlinearity and weak dissipation–if one is interested in the viscosity solution.
@@
In this paper, we use what we call a linear formulation, where we can still approximate the physically relevant solutions – the viscosity solutions – of nonlinear Hamilton-Jacobi equations using a linear representation where the results still remain valid globally in time and truly nonlinear features can be captured. Usually when one computes problems with possible singular solutions–such as shocks in gas dynamics or caustics in Hamilton-Jacobi equations–one needs to use numerical viscosity in order to eliminate numerical oscillations and guarantee convergence to the physically relevant viscosity solutions. Even for a linear $\mathcal{H}$, the numerical viscosity in modern shock or caustic capturing schemes is nonlinear [10]. This adds another difficulty in numerical computation of Hamilton-Jacobi equations. To avoid this numerical difficulty, we use a linear, artificial viscosity approximation, which perturbs the Hamilton-Jacobi equation into a viscous Hamilton-Jacobi equation with a linear viscosity term. The concept of artificial viscosity was first introduced by von Neumann and Ritchmyer [11] in the 1950’s for the computation of shock waves in compressible gas dynamics, and has become a standard numerical tool to handle singular solutions.
To overcome the problem of nonlinear $\mathcal{H}$, we first observe that for quadratic Hamiltonians, the classical Cole–Hopf transformation converts the viscous Hamilton-Jacobi equation into a linear heat equation, providing an exact mapping – thus a true linear representation – between nonlinear and linear dynamics. For more general convex Hamiltonians, we adopt and generalise the entropy penalisation method of Gomes and Valdinoci [1], which constructs an entropy-regularised approximation that correspond to the viscous Hamilton-Jacobi equations. Furthermore, via the Cole-Hopf transform, it gives a linear discrete-time model that corresponds to a linear parabolic PDE, which is convenient for quantum simulation. This transformation enables a linear formulation for the nonlinear viscous Hamilton-Jacobi equation without loss of essential nonlinear behavior. A related method was introduced in [12] to compute the minimal Mather measure.
@@ II.1 Background on viscosity solutions of Hamilton-Jacobi equations
Instead, we use a linear, artificial viscosity approximation. The simplest such approximation is the following regularised problem, called the viscous Hamilton-Jacobi equation
@@
Then by combining Theorem 3.1 and Corollary 3.4 from [23] gives:
@@
$\sup_{0\leq t\leq T}\|\nabla S(t,x)-\nabla S_{\nu}(t,x)\|_{l_{2}(\mathbb{T}^{d})}\leq C{\nu}^{\beta/2},$ (4)
for all $\beta\in(1/2,1)$, $C$ is linear in $d$, scales with $T^{3/2}$, and also depends on $\beta,S_{0}$ and $H$.
@@ II.2.1 Quadratic Hamiltonian
$S_{\nu}(t,x)=-2\nu\ln(\mathcal{N}_{0}u(t,x)),\qquad\mathcal{N}^{2}_{0}=\int e^{-S_{\nu}(0,x)/\nu}dx.$ (6)
@@ II.2.2 More general convex Hamiltonians
In the previous section, the Hamiltonian is quadratic in the canonical momentum, which is represented by the gradient of $S_{\nu}(t,x)$, and intuitively this term acts like a kinetic energy term. For Hamiltonians with terms that are more general in canonical momentum, however, it is not possible to directly transform the nonlinear PDE into a linear one. So we proceed using an alternative method introduced in [1], which still makes use of the Cole-Hopf transformation, though doing so indirectly.
@@
The basic idea proposed in [1] is that, instead of working directly with the PDE in continuous time and then perform the Cole-Hopf transformation, one first chooses a small time step $h>0$ and proceeds with discrete-time schemes. Here we define at time-step $n$ the solution $S_{D}^{n}(x)$ via a time-marching process:
@@
where $G$ is a nonlinear operator that approximately corresponds to the viscous Hamilton-Jacobi equation Eq. (2). The solution of the discrete-time scheme $S_{D}^{n}$ can be related to the solution of another discrete-time scheme $\tilde{u}^{n}$, defined by
$\tilde{u}^{n+1}(x)=\tilde{L}[\tilde{u}^{n}](x)=\int e^{-h\mathcal{L}(x,v)/(2\nu)}\tilde{u}^{n}(x+hv)dv,$ (13)
Here $\tilde{L}$ is a linear operator. Then it has been shown that [1] $S_{D}^{n}(x)$ and $\tilde{u}^{n}$ are related by the Cole-Hopf transformation:
$S_{D}^{n}(x)=-2\nu\ln\tilde{u}^{n}(x).$ (14)
Numerically it will be more convenient to use the normalised equation
${u}^{n+1}(x)={L}[{u}^{n}](x)=\int P(v){u}^{n}(x+hv)dv,\qquad P(v)=\frac{e^{-h{K}(v)/(2\nu)}}{\int e^{-h{K}(v)/(2\nu)}dv},$ (15)
that satisfies the property $\|L[u]\|\leq\|u\|$, hence the dynamics (15) is contractive. It can be shown that (15) approximates some linear parabolic heat-like operator when the discrete time step $h$ is small, as we will see in Lemma 5.
@@
    Solve for $u^{n}(x)$, for $n=0,\cdots,N$, by the linear scheme (15);
@@
    The solution to $S_{\nu}^{N}(x)$ can be recovered by applying the Cole-Hopf transformation:
    $S_{\nu}^{N}(x)=-2\nu\ln u^{N}(x);$ (16)
@@
    $\nabla S_{\nu}^{N}(x)=-2\nu\nabla{u}^{N}(x)/{u}^{N}(x).$ (17)
@@
where an entropy regulariser $\gamma(v)\ln\gamma(v)$ is added for $\nu\neq 0$ and $\gamma(v)$ is a probability density over $\mathbb{R}^{d}$ so $\int_{\mathbb{R}^{d}}\gamma(v)dv=1$. The inclusion of an entropy regularisation [25] introduces uncertainty in $v$ and the optimal probability distribution $\gamma(v)$ is expected to take the form of a Gibbs state with temperature $\nu$, from well-known arguments related to free energy minimisation. Eq. (18) is also a Hopf-Lax-like representation, one that uses an entropy regulariser. Using this definition, one can form a suggestive argument to see how the Hamilton-Jacobi equation emerges. For small $\nu$ and $h$, using Eq. (18), the time derivative of $S_{D}$ behaves roughly like $\partial S_{D}^{n}/\partial t\sim(S_{D}^{n+1}-S_{D}^{n})/h=(G[S_{D}^{n}]-S_{D}^{n})/h\sim\inf(\mathcal{L}-p\cdot v)=-\mathcal{H}(x,p)$, where the $p\cdot v$ term comes from the Taylor expansion $S_{D}^{n}(x+hv)\sim S_{D}^{n}(x)+h\nabla S_{D}^{n}\cdot v$ and $p=-\nabla S_{D}$. This then gives the Hamilton-Jacobi equation $\partial S_{D}/\partial t+\mathcal{H}(x,p)=0$. This is to provide intuition only, and for full details see [1].
@@
The steps in the algorithm above is in discrete-time. However, a continuous-time limit (setting $h\rightarrow 0$) can also be found, which we see later is more suitable for the quantum simulation of PDEs. To derive the continuous-time limit, it is first necessary to ensure that the following linear iterative process is stable
@@
$\frac{\partial u(t,x)}{\partial t}=\sum_{i=1}^{d}\mu_{i}\frac{\partial u(t,x)}{\partial x_{i}}+\sum_{j,k=1}^{d}\nu_{jk}\frac{\partial^{2}u(t,x)}{\partial x_{j}\partial x_{k}},\qquad u(0,x)=\frac{e^{-S(0,x)/(2\nu)}}{\mathcal{N}_{0}},\qquad\mathcal{N}^{2}_{0}=\int e^{-S(0,x)/\nu}dx,$
$\mu_{i}=v_{i}^{*},\qquad\nu_{jk}=v^{*}_{j}v^{*}_{k}+2\nu(D^{-1})_{jk},\qquad v^{*}=\text{argmin}_{v}K(v),\qquad D_{jk}=\partial^{2}K(v^{*})/\partial v_{j}\partial v_{k},$ (26)
where the drift coefficient $\mu_{i}$ and the diffusion coeffients $\nu_{ij}$ are all independent of $x$ and $D$ is a matrix with elements $D_{jk}$. Let $S_{\nu}(t,x)$ be the solution of the viscous Hamilton-Jacobi equation in Eq. (2) with
@@
    Solve for $u(t,x)$ using the linear parabolic PDE in Eq. (5) (using the choice $\mathcal{N}_{0}=1$) in Lemma 5;
@@
    Apply the Cole-Hopf transformation to $u(t,x)$ to obtain $S_{\nu}(t,x)=-2\nu\ln u(t,x)$.
@@ III.2 Digital quantum simulation
If we instead embed the solution of our problem into a discrete-variable quantum state $|u(t)\rangle_{DV}$, then we need to first decompose $x\in[-1/2,1/2]^{d}$ into a grid of size $N_{x}$ in each dimension, and consider the cost of simulating the state of $d\ln N_{x}$ qubits (or a $N_{x}^{d}$-dimensional qudit), that embeds the solution of Eq. (28)
@@
Each $j_{k}$, $k=1,\cdots d$ can be decomposed into its binary decimal representation $j_{k}=j_{k}^{(1)}\cdots j_{k}^{(\ln N_{x})}$, where each $j_{k}^{(l)}\in\{0,1\}$.
@@
which is a matrix of size $N_{\eta}N_{x}^{d}\times N_{\eta}N_{x}^{d}$ that acts on a system of $\ln N_{\eta}+d\ln N_{x}$ qubits. One can obtain $\mathbf{A}_{1,DV}$ and $\mathbf{A}_{2,DV}$ by discretising the spatial components of $\mathbf{A}_{1},\mathbf{A}_{2}$, for example. In this case replace $\hat{p}\rightarrow\hat{P}$, $\hat{x}\rightarrow{X}$, where $\hat{P}_{j}$ is a sparse shift matrix and $\hat{X}=\text{diag}(-1/2N_{x},\cdots,1/2N_{x})$ for the one-dimensional case, and $\hat{D}=\text{diag}(-1/2N_{\eta},\cdots,1/2N_{\eta})$. We summarize the algorithm below.
@@
    Apply projective operator $\hat{\Pi}_{\xi>0}$ on $\ln N_{\xi}$ ancilla qubits (or $N_{\xi}$-dimensional ancilla qudit). Flag success – output $|u(t)\rangle$ – when $\xi>0$ and success probability $p_{succ}=\|\mathbf{u}(t)\|/\|\mathbf{u}(0)\|$ (amplitude amplification used for quadratic boosting).
@@
Let $S_{\nu,\Delta}=-2\nu\ln|u\rangle$ where $|u\rangle$ is the discrete solution obtained by Algorithm 4. Then,
@@
$\tilde{O}\left(\frac{1}{\|\mathbf{u}(t)\|}\left(\alpha_{A}t\mu+\ln\left(\frac{\mu}{\epsilon\|\mathbf{u}(t)\|}\right)\right)\right),\qquad\|\mathbf{u}(0)\|=1,$ (40)
where $\alpha_{A}\geq\|\mathbf{A}_{1}\|,\|\mathbf{A}_{2}\|$ (spectral norm) and using $O(\|\mathbf{u}(0)\|/\|\mathbf{u}(t)\|)$ queries to the initial state preparation protocol for $|v(0)\rangle$. If one discretises the initial ancilla state $|\Xi\rangle_{DV}$ by discretising $\xi$ with the total number of grids $N_{\xi}\sim 1/\epsilon$, then $\mu=O(1/\epsilon)$. There also exists optimal choices of the ancilla initial state that can give $\mu=O(\ln(1/\epsilon))$. For time-dependent Hamiltonians, the optimal query complexity only incurs an extra $\ln(1/\epsilon)$ factor.
@@
Ignoring the terms with logarithmic scaling, one can simulate $|u(t)\rangle_{DV}$ to precision $\epsilon$ with cost $\tilde{O}(\alpha_{A}t\ln(1/\epsilon)/\|\mathbf{u}(t)\|)$, where it is sufficient to choose $\alpha_{A}\geq\max\{\|\mathbf{A}_{1}\|,\|\mathbf{A}_{2}\|\}=O(d^{4}/\epsilon^{2})$.